1) The area of a room is 20m2. If its length is increased by 3m and its width is increased by 1m, the area will be enlarged by 100%. Find the original dimensions of the room.
Answer by JulietG(1812)

L * W = A
L * W = 20m^2
.
If you increase the length by 3 and the width by 1, the area will then be twice as large.
(L+3) * (W+1) = 40m^2
.
Here's one way to do it:
The factors of 40 are -
40,1
20,2
10,4
8,5
The factors of 20 are -
20,1
10,2
5,4
Which of the factors of 20 can be put into the equation (L+3) * (W+1) = 40m^2 to make it true?
Only 5,4 (which becomes 8,5) is possible.
The original dimensions are 5m x 4m. The revised dimensions are 8m x 5m.